Azelastine hydrochloride is an antihistamine nasal spray. A standard size container holds one fluid ounce (oz.) of the liquid. You are searching for this medication in a European drugstore and are asked how many milliliters (mL) there are in one fluid ounce. Using the following conversion factors, determine the number of milliliters in a volume of 1.63 fluid ounces: 1 gallon (gal) = 128 oz., 3.785 x 10-3 cubic meters (m3) = 1 gal, 1 mL = 10-6 m3.
Answer #1

1.63 oz

first, convert to gallon

1.63 oz = 0.012734 gallon

now, convert from gallon to m3

0.012734 gallon = 0.012734 * 3.785e-3 = 4.82e-5 m3

nor, convert from m3 to mL

4.82e-5 m3 = 4.82e-5 / 10-6 = 48.2 mL

therefore,

1 oz = 48.2 mL